OpenStudy (anonymous):

If cos x = .75, what is sin x? 0.87 0.75 0.66 0.69

OpenStudy (anonymous):

okay so you have to solve for x and then plug that value into the sin(x) function

OpenStudy (anonymous):

looks like they want the answer to be positive. sin(x) = sqrt(1 - cos^2(x))

OpenStudy (anonymous):

do sqrt(1 - (0.75)^2)

OpenStudy (anonymous):

so it would be 0.66?

OpenStudy (anonymous):

or would you round?

OpenStudy (anonymous):

actually i think its just a simple arccos question \[\cos(x) = .75\] \[x = \cos^-1(.75)\] and then plug that value into sin(x)

OpenStudy (anonymous):

ohhh ok thanks

OpenStudy (anonymous):

but either way you get the same answer .661

OpenStudy (anonymous):

Thanks alot!